OpenStudy (anonymous):

Can you guys halp me with math pls? Is the sequence 3, 12, 36, ... a geometric sequence? Explain.

OpenStudy (anonymous):

To be in Geometric sequence it must satisfy the following: t_3/t_2 = t_2/t_1

OpenStudy (anonymous):

Does it help?

hartnn (hartnn):

is 12/3 = 36/12 ?

OpenStudy (anonymous):

thanks man

OpenStudy (phi):

geometric sequence multiply to get the next term arithmetic sequence add to get the next term